[figure: 9.7 cm]

ELECTRONIC STRUCTURE AND MAGNETIC PROPERTIES

OF Tb_{3}Rh AND Ho_{3}Rh SINGLE CRYSTALS

E. Talik^{a}, W. Witas^{a}, J. Nowak^{a}, T. Mydlarz^{b}, and A. Winiarski^{a} aInstitute of Physics, University of Silesia, Uniwersytecka 4, 40-007 Katowice, Poland

bInstitute of Low Temperatures and High Fields, Gajowicka 95, 53-529 Wrocław, Poland

Tb_{3}Rh and Ho_{3}Rh belong to the richest in rare earth element compounds in the R − T phase diagrams. They crystallize in the orthorhombic Fe_{3}C - type crystal structure. The Rh atoms are situated in the centre of trigonal prisms, built of Tb and Ho atoms, so they are separated from each other. The formation of a very narrow Rh 4d band was found from the XPS valence band measurements using the electron spectrometer PHI 5700/660. This narrowing of d states in the valence band influences the core level character and magnetic properties of Tb_{3}Rh and Ho_{3}Rh. The X-ray Berg-Barrett topography confirmed the good quality of the single crystals grown by the Czochralski method from a levitated melt. The antiferromagnetic transition was found at 89 K in the ac magnetic susceptibility and in the electrical resistivity temperature dependence for Tb_{3}Rh. Two transitions were observed at 35 K and 25 K in the ac susceptibility and electrical resistivity run for Ho_{3}Rh. The magnetization measurements in the strong magnetic fields are in progress. This work was partly supported by the Foundation for German - Polish Cooperation with the funds from the Federal Republic of Germany.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : E. Talik

Address for correspondence : University of Silesia Institute of Physics Uniwersytecka 4 40-007 Katowice

Email address : talik@us.edu.pl